The base of a ladder is 10 ft away from the wall. The top of the ladder is 11 ft from the floor. Find the length of the ladder to the nearest thousandth.
Answer by checkley77(12569)

10^2+11^2=L^2
100+121=L^2
L^2=221
L=SQRT221
L=14.866 FT. IS THE LENGTH OF THE LADDER.